OpenStudy (anonymous):

Marco is taking a multiple-choice test. There are 5 choices for each question. Marco does not know how to answer one of the questions. If he guesses, what is the probability that he will guess the correct answer?

OpenStudy (anonymous):

p(one correct ans.)=1/5 p(all correct ans.)=\[1/5\times 1/5\times 1/5\times 1/5\times 1/5\]